OpenStudy (anonymous):

The endpoints of a diameter are given. Write an equation of the circle.? (–4, –8) and (1, 0)

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

myininaya (myininaya):

i want this one please

OpenStudy (anonymous):

first if find the midpoint, which is \[(\frac{-4+1}{2},\frac{-8+0}{2})\]

OpenStudy (anonymous):

ok fine

myininaya (myininaya):

\[d=\sqrt{(-4-1)^2+(-8-0)}^2}=\sqrt{25+64}=\sqrt{89\] is the diameter and the radius is \[\sqrt{89\]/2 he already found the center above so the equation is (x-(-3/2))^2+(y-(-4))^2=89/4

myininaya (myininaya):

oops the radius is sqrt{89}/2 it messed up

myininaya (myininaya):

if u can't read the messed up code i just found the length of the diameter using the distance formula